OpenStudy (anonymous):

Each of the numbers 0 through 12 is written on a piece of paper and all of the pieces of paper are placed in a hat. One number is selected at random. Determine the probability that the number selected is even. Note: 0 is considered an even number

OpenStudy (anonymous):

7/13 1/2 7/12 6/13

OpenStudy (asadkarim7):

6/13

OpenStudy (anonymous):

these are choices to chose from

OpenStudy (amistre64):

13 numbers; and 7 evens; 7/13

OpenStudy (anonymous):

you have 13 numbers in the hat. 6 are even. give \[\frac{6}{13}\]

OpenStudy (anonymous):

oops messed up again!

OpenStudy (amistre64):

0,2,4,6,8,10,12

OpenStudy (anonymous):

\[\frac{7}{13}\]

OpenStudy (amistre64):

with great title, comes great responsibility lol

OpenStudy (asadkarim7):

total integers are 13 , you have to select 6 even integers therefore 6/13

OpenStudy (anonymous):

i can't count, but at least i know what \[\sqrt{25}\] is

OpenStudy (anonymous):

so the answer is 7/13? 6/13?

OpenStudy (anonymous):

not the only one who can't count i see 7/13

OpenStudy (asadkarim7):

answer is 6 /13

OpenStudy (anonymous):

ty :)

OpenStudy (amistre64):

13 pieces of paper; 7 marked even, 6 marked odd the chances of getting a paper marked even is 7 out of 13; 7/13

OpenStudy (anonymous):

here is my last question.. its a ven diagram...